OpenStudy (anonymous):

There is a group of boys and girls waiting for school buses. 25 girls get on the bus. The ratio of boys to girls is now 3: 2. 15 boys get on the second bus. There are now the same number of boys and girls at the bus stop. How many students altogether were originally at the bus stop?

OpenStudy (anonymous):

Any help starting this. I know the ratio is 3/5 boys and 2/5 girls at the stop. But blank where to start

OpenStudy (anonymous):

hi @rosegreenr ..shall we start to solve this problem..?

OpenStudy (anonymous):

yes please

OpenStudy (anonymous):

let us start like this.. 3x-15=2x can you make out..how i got that equation

OpenStudy (anonymous):

No. of boys and girls is same after 15 boys taking the second bus. before that there were 3X boys and 2x girls

OpenStudy (anonymous):

@rosegreenr ..do you und stnd that ?

OpenStudy (anonymous):

ok got it so far

OpenStudy (anonymous):

ok ..now 3x-2x=15 x=15 so before 15 boys taking the second bus, no. of boys=3x15=45 no.of girls=2x15=30 so 45:30=3:2

OpenStudy (anonymous):

after 15 boys taking second bus, no. of boyswill be =45-15=30 so no. of girls (30) = no. of boys (30) now

OpenStudy (anonymous):

@rosegreenr are u there?

OpenStudy (anonymous):

slow connection

OpenStudy (anonymous):

clearer now, work backwards

OpenStudy (anonymous):

oh..now read the question once again..and try to sum up no. of students present totally

OpenStudy (mathmath333):

@BIOCRYSTAL whats the final answer coming

OpenStudy (anonymous):

@mathmath333 ..its 100

OpenStudy (anonymous):

@rosegreenr ..got it..?

OpenStudy (mathmath333):

i mean the no of boys and no of girls seperately

OpenStudy (anonymous):

@mathmath333 ..you tell me..how much it would be

OpenStudy (mathmath333):

sry i cant get

OpenStudy (anonymous):

Thanks BIOCRYSTAL, I got the same answer

OpenStudy (anonymous):

@rosegreenr ..can you tell the count separately ..boys and girls

OpenStudy (anonymous):

@rosegreenr and you owe me a medal

OpenStudy (anonymous):

@mathmath333 ..it is 55 girls and 45 boys

OpenStudy (mathmath333):

oh yes i got the same

OpenStudy (anonymous):

@mathmath333 ...ok then go ahead

OpenStudy (mathmath333):

??